OpenStudy (anonymous):

(5+b) x 3/4=9 find b

OpenStudy (kamille):

\[(5+b)*\frac{ 3 }{ 4 }=9|*4\] \[(5+b)*3=36|:3\] 5+b=12 so b is?

OpenStudy (anonymous):

7

OpenStudy (kamille):

correct;)